OpenStudy (anonymous):

Perform the indicated operation : (y4 - 1) ÷ (y + 1)

OpenStudy (dmnts):

\[\frac{ y ^{4}-1 }{ y+1 }=\frac{ (y ^{2}+1)(y ^{2}-1) }{ y+1 }=\frac{ (y ^{2}+1)(y -1)(y+1) }{ y+1 }=(y ^{2}+1)(y -1)=y ^{3}-y ^{2}+y-1\]

OpenStudy (dmnts):

\[(y ^{2}+1)(y-1)=y ^{3}-y ^{2}+y-1\]